# Sentiment Analysis of Financial Text Using Quantum Language Processing QDisCoCirc

### v1

@@ Abstract
We apply quantum distributional compositional circuit (QDisCoCirc) to 3-class sentiment analysis of financial text. In our classical simulations, we keep the Hilbert-space dimension manageable by decomposing each sentence into short contiguous chunks. Each chunk is mapped to a shallow quantum circuit, and the resulting Bloch vectors are used as a sequence of quantum tokens. Simple averaging of chunk vectors ignores word order and syntactic roles. We therefore add a small Transformer encoder over the raw Bloch-vector sequence and attach a CCG-based type embedding to each chunk. This hybrid design preserves physically interpretable semantic axes of quantum tokens while allowing the classical side to model word order and long-range dependencies. The sequence model improves test macro-F1 over the averaging baseline and chunk-level attribution further shows that evidential mass concentrates on a small number of chunks, that type embeddings are used more reliably for correctly predicted sentences. For real-world quantum language processing applications in finance, future key challenges include circuit designs that avoid chunking and the design of inter-chunk fusion layers.
@@ 1 Introduction
This paper investigates how QDisCoCirc enables causal and intervention-based analysis in quantum language processing. Combinatory Categorial Grammar (CCG) is a framework that systematically composes word-meaning representations according to grammatical structure. Laakkonen et al. [9] extended DisCoCirc, which generalizes CCG to text circuits, to quantum computation and proposed efficient algorithms for constructive text processing; they referred to the resulting framework as QDisCoCirc. QDisCoCirc performs compositional semantics using sequences of quantum gates and provides explicit interpretability by identifying, gate by gate, which word influences which qubit. Duneau et al. [5] implemented QDisCoCirc on Quantinuum’s H1-1 trapped-ion quantum processor. By reusing qubits under the constraints of the 20-qubit H1-1 device, they managed to fit circuits of up to 108 logical qubits into just 20 physical qubits. They confirmed that there was no significant loss of accuracy even for texts containing up to 30 entities (noun phrases), demonstrating the scalability of their approach to longer texts.
In this study, we apply QDisCoCirc to sentiment analysis of financial texts and evaluate its performance via classical simulation. The task is sentiment classification with three classes-negative, neutral, and positive-extending the binary QNLP experiments of Duneau et al. [5]. By applying QDisCoCirc to realistic financial sentiment tasks, we provide, to the best of our knowledge, the first case study of quantum language processing in real-world finance. While Duneau et al. [5] discuss inspectability only at the circuit-component (word state/box) level, our study differs in that it allows us to decompose, within a sentence, the attribution of “which information channels contributed and to what extent” into three components: (a) Bloch-vector representations (semantic axes), (b) CCG types (syntactic channel), and (c) type gates (control of throughput in the syntactic channel). In this way, our study complements the hardware-based demonstration of scalability by Duneau et al. [5] and provides a novel evaluation perspective that enables quantitative tracking of the respective roles of quantum (Bloch) and syntactic (CCG) factors at the sentence level.
@@ 2 Model
QDisCoCirc is a framework that represents sentences as quantum circuits built from sequential and parallel compositions of small-scale components. Content words such as nouns are represented as input states (“states”); verbs and adjectives are processes (“boxes”) that transform or combine states; and sentence evaluation is represented as an effect. Sentences can also be connected via circuit composition, and semantic similarity between sentences is evaluated by the overlap of the output states of the corresponding circuits. Grammar serves as a wiring specification that determines which wires are contracted and which remain.
@@ 3.1 Normalization of Input Representations via Vocabulary Rewriting
    Preventing structural errors during circuit construction
    When decomposing a sentence into semantic units to form a circuit, even a slight misalignment at the “connection points” between words or phrases can prevent the circuit from being constructed properly. To avoid this, words with the same grammatical role are normalized to a common tag, ensuring that connections still align correctly after chunking.
    Example: Frequently appearing nouns such as earnings, dividend, and asset price are mapped to tags representing “financial indicators,” whereas verbs such as cuts, raises, and misses are mapped to tags indicating upward or downward movement. This allows similar vocabulary to be treated as the same component (see Appendix A for detailed tagging procedures).
@@
    Handling unknown and rare words
    Here, “unknown and rare words” refer to words that appear at test but were rare or absent in the training data. First, during chunking, syntactic analysis based on CCG assigns each word a type, such as noun type or verb type, allowing the syntactic role of even out-of-vocabulary words to be identified. Furthermore, if a word belongs to one of the financial vocabulary clusters defined during preprocessing, it is mapped to the corresponding tag. If there is no applicable tag, the word is assigned its own rotation-angle parameter, which remains at its initial value. In this way, even when unknown words appear at test, the structure of the quantum circuit remains intact and inference can proceed.
@@ 3.2 Preprocessing to Preserve Chunk Boundaries
    Passive constructions and fixed expressions
    Passive constructions such as is expected to and was announced to, as well as fixed expressions frequently used in financial news (e.g., is set to), are treated as single chunks. This standardizes the connection points of verb phrases and stabilizes the links between chunks.
@@
    Compound prepositions
    Multiword prepositions such as due to and as a result of must be treated as single units; otherwise, their meaning is fragmented. Treating them as single chunks makes the corresponding prepositional phrases well-formed units for composition.
@@
    Proper nouns such as company names
    We treat expressions such as Company A Inc. and Company A as a single chunk.
@@
    URLs and symbol sequences
    URLs and special symbol sequences are long character strings that break the assumed syntactic structure for circuit construction. Replacing them with single-word chunks prevents such structural disruptions.
@@
    as + past participle (+ adverb)
    Subordinate clauses such as as expected and as widely are attached to the subsequent main-clause chunk, preventing these clauses from being split during chunking.
@@
    to + verb
    Infinitival phrases (to + verb) form separate units if left unmodified. We attach them to their subjects or governing verbs to keep the resulting chunk structure compositionally continuous.
@@
    Discourse markers at sentence start
    Sentence-initial discourse markers such as However, and Therefore, do not directly contribute to the core semantic structure. If left as independent chunks, they introduce redundant nodes or lead to type mismatches in the circuit. To prevent this, we do not isolate them as separate chunks but instead combine them with the following main clause, thereby avoiding an increase in spurious nodes.
@@ 4 Experiment
QDisCoCirc 0.696 0.551 $2.2\times 10^{4}$
@@ 5 Confidence and Sentence-Level Explanation Metrics
In this section, we introduce two types of scores to enhance the interpretability of the prediction results: (A) probability-based confidence metrics and (B) sentence-level sensitivity metrics based on representation-level interventions. Metrics in group (A) are conventional probability-based scores, whereas metrics in group (B) measure the sensitivity of the output when sentence-level diagram surgery [12] is applied to the quantum state representations (Bloch components) of QDisCoCirc.
@@ 7 Conclusion
This study applied QDisCoCirc to three-class sentiment classification of financial text and evaluated the effectiveness of quantum compositional models on realistic financial tasks. By augmenting QDisCoCirc with a shallow Transformer encoder that models sentence structure, we showed that it is possible to partially overcome the limitations of mean pooling. Moving forward, it will be important to balance performance and scalability by exploiting the unique features of QDisCoCirc, developing new compositional rules that combine information between chunks at the quantum level, and incorporating circuit-compression techniques such as XZ reduction. Implementing and evaluating financial sentiment analysis tasks on actual devices, such as superconducting quantum processors, is another key next step.
@@ 1. Lexical Level: Normalizing Words
    Unifying copulas (be-verbs)
    Map is, are, was, etc. to a single tag.
@@
    Normalization of numbers and units
    Normalize terms such as million, percent, USD, and kWh into numeric/unit tags.
@@
    Clustering upward/downward movements
    Group upward-movement verbs such as rise, increase, surge and downward-movement verbs such as fall, decline into corresponding tags.
@@ 2. Phrase Level: Normalizing Meaning
    Tagging finance-specific relative pronouns
    Tag words such as “that”, “which”, and “where” in a way that reflects their use in financial news.
@@
    Combining compound prepositions
    Merge multiword prepositional phrases such as “due to” and “as a result of” into unified tags.
@@ 3. Syntactic Level: Normalizing Function Words
    Unifying comparative expressions
    Standardize words and phrases such as “higher”, “lower”, “better”, “worse”, and “compared”.

### v2

# Diagram-to-Circuit QNLP for Financial Sentiment Analysis

@@ Abstract
We study a QDisCoCirc-inspired, chunked diagram-to-circuit quantum natural language processing (QNLP) model for three-class sentiment classification of financial texts. In our classical simulations, we keep the Hilbert-space dimension manageable by decomposing each sentence into short contiguous chunks. Each chunk is mapped to a shallow quantum circuit, and the resulting Bloch vectors are used as a sequence of quantum tokens. Simple averaging of chunk vectors ignores word order and syntactic roles. We therefore add a small Transformer encoder over the raw Bloch-vector sequence and attach a CCG-based type embedding to each chunk. This hybrid design preserves physically interpretable semantic axes of quantum tokens while allowing the classical side to model word order and long-range dependencies. The sequence model improves test macro-F1 over the averaging baseline and chunk-level attribution further shows that evidential mass concentrates on a small number of chunks, that type embeddings are used more reliably for correctly predicted sentences. For real-world quantum language processing applications in finance, future key challenges include circuit designs that avoid chunking and the design of inter-chunk fusion layers.
@@ 1 Introduction
This paper investigates how a QDisCoCirc-inspired, chunked diagram-to-circuit model enables causal and intervention-based analysis in quantum natural language processing. Combinatory Categorial Grammar (CCG) is a framework that systematically composes word-meaning representations according to grammatical structure. Laakkonen et al. [9] extended DisCoCirc, which generalizes CCG to text circuits, to quantum computation and proposed efficient algorithms for constructive text processing; they referred to the resulting framework as QDisCoCirc. QDisCoCirc performs compositional semantics using sequences of quantum gates and provides explicit interpretability by identifying, gate by gate, which word influences which qubit. Duneau et al. [5] implemented QDisCoCirc on Quantinuum’s H1-1 trapped-ion quantum processor. By reusing qubits under the constraints of the 20-qubit H1-1 device, they managed to fit circuits of up to 108 logical qubits into just 20 physical qubits. They confirmed that there was no significant loss of accuracy even for texts containing up to 30 entities (noun phrases), demonstrating the scalability of their approach to longer texts.
In this study, we evaluate a QDisCoCirc-inspired, chunked diagram-to-circuit QNLP model on sentiment analysis of financial texts via classical simulation. The task is sentiment classification with three classes-negative, neutral, and positive-extending the binary QNLP experiments of Duneau et al. [5]. By applying this QDisCoCirc-inspired approach to realistic financial sentiment tasks, we provide a finance-oriented case study and introduce intervention-based explanation metrics that quantify the roles of the quantum and classical components. While Duneau et al. [5] discuss inspectability only at the circuit-component (word state/box) level, our study differs in that it allows us to decompose, within a sentence, the attribution of “which information channels contributed and to what extent” into three components: (a) Bloch-vector representations (semantic axes), (b) CCG types (syntactic channel), and (c) type gates (control of throughput in the syntactic channel). In this way, our study complements the hardware-based demonstration of scalability by Duneau et al. [5] and provides a novel evaluation perspective that enables quantitative tracking of the respective roles of quantum (Bloch) and syntactic (CCG) factors at the sentence level.
@@
Terminology and positioning.
The QDisCoCirc framework [5] builds on the DisCoCirc “text-as-circuit” view and aims at compiling entire sentences (and, in principle, longer texts) into executable circuits. In this paper, due to classical-simulation constraints, we do not attempt to compile each whole sentence into a single end-to-end text circuit. Instead, we instantiate compositional circuits at the level of short CCG-derived chunks: each chunk is converted into a DisCoCat-style pregroup diagram and mapped to a shallow quantum circuit, while inter-chunk context is handled classically (pooling or a small Transformer). We therefore use the term QDisCoCirc-inspired when referring to our instantiation.
@@ 2 Model
QDisCoCirc is a framework that represents sentences as quantum circuits built from sequential and parallel compositions of small-scale components. Due to the limits of classical simulation, our implementation constructs shallow circuits only at the level of short CCG-derived chunks (via DisCoCat-style pregroup diagrams) and models inter-chunk context on the classical side. Content words such as nouns are represented as input states (“states”); verbs and adjectives are processes (“boxes”) that transform or combine states; and sentence evaluation is represented as an effect. Sentences can also be connected via circuit composition, and semantic similarity between sentences is evaluated by the overlap of the output states of the corresponding circuits. Grammar serves as a wiring specification that determines which wires are contracted and which remain.
@@ 3.1 Normalization of Input Representations via Vocabulary Rewriting
    Preventing structural errors during circuit constructionWhen decomposing a sentence into semantic units to form a circuit, even a slight misalignment at the “connection points” between words or phrases can prevent the circuit from being constructed properly. To avoid this, words with the same grammatical role are normalized to a common tag, ensuring that connections still align correctly after chunking.Example: Frequently appearing nouns such as earnings, dividend, and asset price are mapped to tags representing “financial indicators,” whereas verbs such as cuts, raises, and misses are mapped to tags indicating upward or downward movement. This allows similar vocabulary to be treated as the same component (see Appendix A for detailed tagging procedures).
@@
    Handling unknown and rare wordsHere, “unknown and rare words” refer to words that appear at test but were rare or absent in the training data. First, during chunking, syntactic analysis based on CCG assigns each word a type, such as noun type or verb type, allowing the syntactic role of even out-of-vocabulary words to be identified. Furthermore, if a word belongs to one of the financial vocabulary clusters defined during preprocessing, it is mapped to the corresponding tag. If there is no applicable tag, the word is assigned its own rotation-angle parameter, which remains at its initial value. In this way, even when unknown words appear at test, the structure of the quantum circuit remains intact and inference can proceed.
@@ 3.2 Preprocessing to Preserve Chunk Boundaries
    Passive constructions and fixed expressionsPassive constructions such as is expected to and was announced to, as well as fixed expressions frequently used in financial news (e.g., is set to), are treated as single chunks. This standardizes the connection points of verb phrases and stabilizes the links between chunks.
@@
    Compound prepositionsMultiword prepositions such as due to and as a result of must be treated as single units; otherwise, their meaning is fragmented. Treating them as single chunks makes the corresponding prepositional phrases well-formed units for composition.
@@
    Proper nouns such as company namesWe treat expressions such as Company A Inc. and Company A as a single chunk.
@@
    URLs and symbol sequencesURLs and special symbol sequences are long character strings that break the assumed syntactic structure for circuit construction. Replacing them with single-word chunks prevents such structural disruptions.
@@
    as + past participle (+ adverb)Subordinate clauses such as as expected and as widely are attached to the subsequent main-clause chunk, preventing these clauses from being split during chunking.
@@
    to + verbInfinitival phrases (to + verb) form separate units if left unmodified. We attach them to their subjects or governing verbs to keep the resulting chunk structure compositionally continuous.
@@
    Discourse markers at sentence startSentence-initial discourse markers such as However, and Therefore, do not directly contribute to the core semantic structure. If left as independent chunks, they introduce redundant nodes or lead to type mismatches in the circuit. To prevent this, we do not isolate them as separate chunks but instead combine them with the following main clause, thereby avoiding an increase in spurious nodes.
@@ 4 Experiment
QDisCoCirc-inspired (chunked) 0.696 0.551 $2.2\times 10^{4}$
@@ 5 Confidence and Sentence-Level Explanation Metrics
In this section, we introduce two types of scores to enhance the interpretability of the prediction results: (A) probability-based confidence metrics and (B) sentence-level sensitivity metrics based on representation-level interventions. Metrics in group (A) are conventional probability-based scores, whereas metrics in group (B) measure the sensitivity of the output when sentence-level diagram surgery [12] is applied to the quantum state representations (Bloch components) of our QDisCoCirc-inspired chunk circuits.
@@ 7 Conclusion
This study evaluated a QDisCoCirc-inspired, chunked diagram-to-circuit QNLP model for three-class sentiment classification of financial texts in a classical-simulation setting. By augmenting QDisCoCirc with a shallow Transformer encoder that models sentence structure, we showed that it is possible to partially overcome the limitations of mean pooling. Moving forward, it will be important to balance performance and scalability by exploiting the unique features of QDisCoCirc, developing new compositional rules that combine information between chunks at the quantum level, and incorporating circuit-compression techniques such as XZ reduction. Implementing and evaluating financial sentiment analysis tasks on actual devices, such as superconducting quantum processors, is another key next step.
@@
Acknowledgements
We thank Konstantinos Meichanetzidis for helpful discussions and feedback.
@@ 1. Lexical Level: Normalizing Words
    Unifying copulas (be-verbs)Map is, are, was, etc. to a single tag.
@@
    Normalization of numbers and unitsNormalize terms such as million, percent, USD, and kWh into numeric/unit tags.
@@
    Clustering upward/downward movementsGroup upward-movement verbs such as rise, increase, surge and downward-movement verbs such as fall, decline into corresponding tags.
@@ 2. Phrase Level: Normalizing Meaning
    Tagging finance-specific relative pronounsTag words such as “that”, “which”, and “where” in a way that reflects their use in financial news.
@@
    Combining compound prepositionsMerge multiword prepositional phrases such as “due to” and “as a result of” into unified tags.
@@ 3. Syntactic Level: Normalizing Function Words
    Unifying comparative expressionsStandardize words and phrases such as “higher”, “lower”, “better”, “worse”, and “compared”.